OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x + 4)(2x – 1)(x – 7)?

OpenStudy (luigi0210):

well it's already factored, just set each one equal to zero: x+4=0 2x-1=0 x-7=0

OpenStudy (anonymous):

Each factor needs to be zero. So that means: x - 4 = 0 x = 4 2x - 1 = 0 2x = 1 x = 1/2 x + 7 = 0 x = -7 Does that help?

OpenStudy (anonymous):

okkk so its -4,1/2,and 7 ?

OpenStudy (luigi0210):

yup :) MK messed up

OpenStudy (anonymous):

Crap I did just noticed sorry @Lesha

OpenStudy (anonymous):

its okk (: . you still helped me understand the concept Which of the following represent the zeros of f(x) = 6x3 – 31x2 + 4x + 5 ? how do i find this one ?

OpenStudy (luigi0210):

mistakes happen ^_^'

OpenStudy (anonymous):

I think f(5) = 6*125 - 31*25 + 4*5 + 5 = 750 - 775 + 20 + 5 = 0 f(-1/3) = -6/27 - 31/9 - 4/3 + 5 = -2/9 - 31/9 - 4/3 + 5 = 0 f(1/2) = 6/8 - 31/4 + 4/2 + 5 = 3/4 - 31/4 + 2 + 5 = 0 So the zeros are 5, -1/3, and 1/2.

OpenStudy (anonymous):

Hope thats right

OpenStudy (luigi0210):

How did you get those? @MKTY

OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 2x2 – 24x ? –6, 0, 4

OpenStudy (anonymous):

is that right ?

OpenStudy (luigi0210):

Yes, that's right :P

OpenStudy (anonymous):

(;;;; What are the possible number of positive, negative, and complex zeros of f(x) = 3x4 – 5x3 – x2 – 8x + 4 ?

ganeshie8 (ganeshie8):

sign changes

ganeshie8 (ganeshie8):

do step by step

OpenStudy (anonymous):

2

ganeshie8 (ganeshie8):

step by step

ganeshie8 (ganeshie8):

step1 : f(x) sign changes step2 : f(-x) sign changes

OpenStudy (anonymous):

2 then 3 ?

ganeshie8 (ganeshie8):

2 for f(x) is correct, so that gives a positive zeroes of : 2 or 0

ganeshie8 (ganeshie8):

3 for f(-x) is wrong, try again

ganeshie8 (ganeshie8):

f(x) = = 3x4 – 5x3 – x2 – 8x + 4 f(-x) = 3x4 + 5x3 - x2 + 8x + 4 how many sign changes ?

OpenStudy (anonymous):

ummm 2. how do you know what to chhange it to ?

ganeshie8 (ganeshie8):

good q :) lets work on it a bit

ganeshie8 (ganeshie8):

f(x) = = 3x4 – 5x3 – x2 – 8x + 4 for f(-x), u just replace x with -x everywhere

ganeshie8 (ganeshie8):

replace and see wat u get

OpenStudy (anonymous):

like this 1f(x) = = -3x4 – 5x3 +x2 – 8x - 4 ? soorryy im confused

ganeshie8 (ganeshie8):

not exactly, just replace x with -x first :- \(3x^4 – 5x^3 – x^2 – 8x + 4 \) \(3(-x)^4 – 5(-x)^3 – (-x)^2 – 8(-x) + 4 \)

OpenStudy (anonymous):

okk i get it. so now whats the next step.

ganeshie8 (ganeshie8):

u get how to figure out f(-x) ?

OpenStudy (anonymous):

yess i dooo nowww.

ganeshie8 (ganeshie8):

great !

ganeshie8 (ganeshie8):

once we knw number of sign changes in f(x) and f(-x), here is how we conclude :-

ganeshie8 (ganeshie8):

2 sign changes in f(x) . so number of positive zeroes = 2, or 0 2 sign changes in f(-x) . so number of negative zeroes = 2, or 0 Number of complex zeroes : 4, 2, or 0

OpenStudy (anonymous):

According to The Fundamental Theorem of Algebra, how many zeros does the function f(x) = 15x23 + 41x19 + 13x5– 10 have ?

ganeshie8 (ganeshie8):

what does "Fundamental Theorem of Algebra" say ?